# Mitigating Negative Flips via Margin Preserving Training

### v1

@@ Introduction
Negative flips in classification tasks are most likely to occur when the number of classes increases to accommodate new categories (Yan et al. 2021; Zhao et al. 2024; Zhang et al. 2021). From a margin-based perspective, increasing the number of classes at each model update reduces the margin between adjacent classes, due to class means pushing themselves to arrange in a way that maximizes the minimum one-vs-rest margin (Jiang et al. 2024a; Papyan, Han, and Donoho 2020; Yaras et al. 2022; Tirer, Huang, and Niles-Weed 2023). In addition to the reduction imposed by the optimization process and geometric constraints (Jiang et al. 2024a), when new categories are introduced, the decision margins are subject to further challenges arising from negative transfer (Li, Nguyen, and Zhang 2023; Zhang et al. 2022). As the model learns to encode features that discriminate among an expanded set of classes, the incorporation of additional categories may lead to the emergence of interfering features, thereby adversely impacting the performance on previously acquired tasks or classes (Tirer, Huang, and Niles-Weed 2023; Liu et al. 2021; Torralba, Murphy, and Freeman 2007). In extreme cases, negative flips involve not only ambiguous samples near the decision boundaries but also high-confidence samples influenced, for example, by spurious correlations (Wang et al. 2021).
@@
[…]
@@
For CIFAR100 and CIFAR-10, we train a ResNet-18 and a ResNet-50 (He et al. 2016) using SGD (momentum = 0.9) with a base learning rate of 0.1, applying cosine annealing (Loshchilov and Hutter 2017) over 100 epochs with a batch size of 128. For ImageNet1k, we adopt the standard ImageNet1k training recipe provided directly by PyTorch repository. Specifically, we train a ResNet-18 and a ResNet-50 (He et al. 2016) with embedding dimension 128 using SGD (momentum = 0.9) with an initial learning rate of 0.1, decayed by a factor of ten every 30 epochs for a total of 90 epochs. The batch size is set to 128. For our method, we set $\lambda=0.4$ and $\lambda=1.0$ when LM of Equation 6 and KL of Equation 5 are used as distillation functions, respectively. Following prior work (Yan et al. 2021; Zhao et al. 2024), we train every model from scratch, as this represents the most challenging setting, resulting in the highest NFR (Yan et al. 2021).
Figure 3: Comparison of embedding spaces for ResNet-18 models (embedding size = 2) on the CIFAR-10 test set under various update strategies. Notably, when the model is trained directly on multiple classes (“No Treatment”), the inter-class margins are narrower compared to the original model. Our proposed Margin Preservation Training (MPT) maintains the margins of the original model, thereby significantly reducing the rate of negative flips without compromising the error rate across all classes.
@@
To complement the quantitative analysis presented in previus section, we provide qualitative results obtained from training a ResNet-18 model with a two-dimensional feature space on CIFAR-10 using different methods. A two-dimensional feature space enables direct visualization of the learned representations without the need for dimensionality reduction techniques such as t-SNE (Van der Maaten and Hinton 2008) or UMAP (McInnes, Healy, and Melville 2018). Figure 3 presents the resulting representations across six different settings: (1) “Old Model”, (2) “No Treatment”, (3) PCT-KL, (4) MPT (no bias, $k=0$), (5) MPT (no distillation, $\lambda=0$ in Equation 3), and (6) MPT. In each subplot, data points are color-coded according to their true class labels.
In the "Old Model" each class forms a compact cluster centered within its respective angular sector and is well-separated from adjacent decision boundaries. In "No Treatment" these clusters drift outward and increasingly intersect neighboring sectors; specifically, the green and blue points cross their boundaries, as do the red and newly introduced brown classes. This behavior evidences significant margin shrinkage and a tighter geometric arrangement compared to the old model, consistent with the findings of (Wu et al. 2022). The PCT-KL method attempts to replicate the feature distribution of the Old Model, resulting in clusters that are more concentrated near the origin compared to the "No Treatment" scenario. However, not all old class margins are well preserved: some red class points are within the brown class region, indicating a uniform reduction of margins for that class. MPT (no bias, $k=0$) produces effects similar to those of PCT-KL. By leveraging knowledge from the new model, it prevents the collapse of new class clusters toward the center or into regions associated with old classes, resulting in a slightly lower error rate. MPT (no distillation) clearly preserves the margins of the old classes, as all old class points remain within their respective regions. However, this approach results in a high error rate for the new classes, as they are mapped into regions corresponding to the old classes, thereby yielding large decision margins for the new classes. Our proposed approach, MPT, achieves the best results. It provides superior margin preservation for old classes and a lower error rate for new classes, with each point correctly residing in its respective class region. Figure 3 shows the emergence of interfering features with the introduction of new categories, as discussed by Tirer, Huang, and Niles-Weed (2023). For instance, the new brown class (Dog) leads to interference in the learning of the old red class (Cat), as they are semantically similar.
@@
To better understand the contributions of each component in our proposed method, we analize the effects of margin-based regularization and double-source focal distillation. Table 3 illustrate the impact of removing margin-based regularization ($k=0$ in Equation 1) or double-source focal distillation by setting $\lambda=0$ in Equation 3. The results demonstrate that both margin-based regularization and focal distillation are essential for minimizing negative flips. Removing either component increases the Negative Flip Rate (NFR) and Relative Negative Flip Rate (Rel-NFR), highlighting their complementary effects. Margin-based regularization alone improves performance on old classes but reduces accuracy on new classes, while distillation alone increases negative flips. Our proposed methods, MPT-KL and MPT-LM, which combine both components, consistently achieve superior performance compared to variants using only one. This confirms that integrating margin preservation and focal distillation is crucial for balanced performance, minimizing negative flips, and maintaining high accuracy across both old and new classes.

### v2

@@ Introduction
Negative flips in classification tasks are most likely to occur when the number of classes increases to accommodate new categories (Yan et al. 2021; Zhao et al. 2024; Zhang et al. 2021). From a margin-based perspective, increasing the number of classes at each model update reduces the margin between adjacent classes, due to class means pushing themselves to arrange in a way that maximizes the minimum one-vs-rest margin (Jiang et al. 2024a; Papyan, Han, and Donoho 2020; Yaras et al. 2022; Tirer, Huang, and Niles-Weed 2023; Pernici et al. 2019, 2021). In addition to the reduction imposed by the optimization process and geometric constraints (Jiang et al. 2024a), when new categories are introduced, the decision margins are subject to further challenges arising from negative transfer (Li, Nguyen, and Zhang 2023; Zhang et al. 2022). As the model learns to encode features that discriminate among an expanded set of classes, the incorporation of additional categories may lead to the emergence of interfering features, thereby adversely impacting the performance on previously acquired tasks or classes (Tirer, Huang, and Niles-Weed 2023; Liu et al. 2021; Torralba, Murphy, and Freeman 2007). In extreme cases, negative flips involve not only ambiguous samples near the decision boundaries but also high-confidence samples influenced, for example, by spurious correlations (Wang et al. 2021).
@@
Figure 2: Schematic overview of the proposed Margin Preserving Training (MPT) approach. A model $\phi$ is trained with the margin-calibrated loss $\mathcal{L}_{\Delta}$ while receiving double-source focal distillation from two reference models, $\phi_{\text{old}}$ and $\phi_{\text{new}}$ (both optimized with cross-entropy). This combination mitigates margin underestimation for new classes, reduces negative flips between model updates, and preserves decision margins for old classes. The gray region marks areas near decision boundaries where negative flips typically occur. Margins are illustrated in a 2D embedding space for clarity, although MPT operates on logit-space margins over all classes.
@@
For CIFAR100 and CIFAR10, we train a ResNet-18 and a ResNet-50 (He et al. 2016) using SGD (momentum = 0.9) with a base learning rate of 0.1, applying cosine annealing (Loshchilov and Hutter 2017) over 100 epochs with a batch size of 128. For ImageNet1k, we adopt the standard ImageNet1k training recipe provided directly by PyTorch repository. Specifically, we train a ResNet-18 and a ResNet-50 (He et al. 2016) with embedding dimension 128 using SGD (momentum = 0.9) with an initial learning rate of 0.1, decayed by a factor of ten every 30 epochs for a total of 90 epochs. The batch size is set to 128. For our method, we set $\lambda=0.4$ and $\lambda=1.0$ when LM of Equation 6 and KL of Equation 5 are used as distillation functions, respectively. Following prior work (Yan et al. 2021; Zhao et al. 2024), we train every model from scratch, as this represents the most challenging setting, resulting in the highest NFR (Yan et al. 2021).
Figure 3: Comparison of embedding spaces for ResNet-18 models (embedding size = 2) on the CIFAR10 test set under various update strategies. Notably, when the model is trained directly on multiple classes (“No Treatment”), the inter-class margins are narrower compared to the original model. Our proposed Margin Preservation Training (MPT) maintains the margins of the original model, thereby significantly reducing the rate of negative flips without compromising the error rate across all classes.
@@
To complement the quantitative analysis presented in previus section, we provide qualitative results obtained from training a ResNet-18 model with a two-dimensional feature space on CIFAR10 using different methods. A two-dimensional feature space enables direct visualization of the learned representations without the need for dimensionality reduction techniques such as t-SNE (Van der Maaten and Hinton 2008) or UMAP (McInnes, Healy, and Melville 2018). Figure 3 presents the resulting representations across six different settings: (1) “Old Model”, (2) “No Treatment”, (3) PCT-KL, (4) MPT (no bias, $k=0$), (5) MPT (no distillation, $\lambda=0$ in Equation 3), and (6) MPT. In each subplot, data points are color-coded according to their true class labels.
In the "Old Model" each class forms a compact cluster centered in its angular sector and is well-separated from adjacent decision boundaries. In "No Treatment" these clusters drift outward and increasingly intersect neighboring sectors: the green and blue points cross their boundaries, as do the red and newly introduced brown classes. This indicates substantial margin shrinkage and a tighter geometric arrangement relative to the old model, consistent with the findings of (Wu et al. 2022). The PCT-KL method attempts to replicate the feature distribution of the Old Model, resulting in clusters that are more concentrated near the origin compared to the "No Treatment" scenario. However, not all old class margins are well preserved: some red class points are within the brown class region, indicating a uniform reduction of margins for that class. MPT (no bias, $k=0$) produces effects similar to those of PCT-KL. By leveraging knowledge from the new model, it prevents new class clusters from collapsing toward the center or into old class regions, resulting in a slightly lower error rate. MPT (no distillation) clearly preserves the margins of the old classes, as all old class points remain within their respective regions. However, this approach results in a high error rate for the new classes, as they are mapped into regions corresponding to the old classes, thereby yielding large decision margins for the new classes. Our proposed approach, MPT, achieves the best results. It provides superior margin preservation for old classes and a lower error rate for new classes, with each point correctly residing in its respective class region. Figure 3 shows the emergence of interfering features with the introduction of new categories, as discussed by Tirer, Huang, and Niles-Weed (2023). For instance, the new brown class (Dog) leads to interference in the learning of the old red class (Cat), as they are semantically similar.
@@
To better understand the contributions of each component in our method, we analize the effects of margin-based regularization and double-source focal distillation. Table 3 shows the impact of removing margin-based regularization ($k=0$ in Equation 1) or double-source focal distillation by setting $\lambda=0$ in Equation 3. The results show that both margin-based regularization and focal distillation are essential for minimizing negative flips. Removing either component increases the Negative Flip Rate (NFR) and Relative Negative Flip Rate (Rel-NFR), highlighting their complementary effects. Margin-based regularization alone improves performance on old classes but reduces accuracy on new classes, while distillation alone increases negative flips. Our proposed methods, MPT-KL and MPT-LM, which combine both components, consistently achieve superior performance compared to variants using only one. This confirms that integrating margin preservation and focal distillation is crucial for balanced performance, minimizing negative flips, and maintaining high accuracy across both old and new classes.